### v1

# Rethinking Causal Discovery Through the Lens of Exchangeability

Abstract
Causal discovery methods have traditionally been developed under two distinct regimes: independent and identically distributed (i.i.d.) and timeseries data, each governed by separate modelling assumptions. In this paper, we argue that the i.i.d. setting can and should be reframed in terms of exchangeability, a strictly more general symmetry principle. We present the implications of this reframing, alongside two core arguments: (1) a conceptual argument, based on extending the dependency of experimental causal inference on exchangeability to causal discovery; and (2) an empirical argument, showing that many existing i.i.d. causal‐discovery methods are predicated on exchangeability assumptions, and that the sole extensive widely-used real‐world “i.i.d.” benchmark (the Tübingen dataset) consists mainly of exchangeable (and not i.i.d.) examples. Building on this insight, we introduce a novel synthetic dataset [^1] All our code is open-source and available in tiago.brogueira@github.com that enforces only the exchangeability assumption, without imposing the stronger i.i.d. assumption. We show that our exchangeable synthetic dataset mirrors the statistical structure of the real‐world ”i.i.d.” dataset more closely than all other i.i.d. synthetic datasets. Furthermore, we demonstrate the predictive capability of this dataset by proposing a neural‐network–based causal‐discovery algorithm trained exclusively on our synthetic dataset, and which performs similarly to other state‐of‐the‐art i.i.d. methods on the real‐world benchmark.
Keywords: cause-effect pairs, synthetic dataset, causal inference, neural network, statistical learning
1 Introduction
The aim of scientific research is often to find causal relationships between certain variables of interest. This stems either from a desire to intervene in the system under study, (and not merely be able to make accurate statements over the data we observe) or from a goal of uncovering causal mechanisms underlying the observations, seeking a deeper understanding of the underlying phenomena (Pearl, 2009). Traditionally, these relationships are uncovered by performing experiments (known in this context as interventions). However, this can often be impossible or impractical (Glymour et al., 2019). Furthermore, if not all intervened variables are controlled for, this can lead to erroneous results (e.g., Simpson’s paradox (Ameringer et al., 2009)). In these situations, the need arises to learn causal relationships from just observational data. This growing field, which uses a purely data-driven approach to learn causal structure (represented using a graph), is called causal discovery.
Causal discovery methods are typically divided into two main categories: those for independent and identically distributed (i.i.d.) data and those for time-series data (Hasan et al., 2023), as the latter inherently violate the i.i.d. assumption underlying static causal models (Günther et al., 2023). This work will not address time-series data. [^2] From this point onwards, we will often refer simply to causal discovery, but it should be interpreted as referring to the family of causal discovery focused on i.i.d. data. An important result in this field is that, when working with i.i.d. purely observational data and without further assumptions, one can only identify the causal structure up to the so-called Markov equivalence class (MEC), which is defined as the set of causal graphs that satisfy the same conditional independence properties (Spirtes et al., 2001). In particular, in the case of a pair of dependent variables, $X$ and $Y$, both causal directions ($X\rightarrow Y$, $Y\rightarrow X$) belong to the same MEC, as there are no conditional independence properties to be satisfied. Furthermore, with i.i.d. observations, everything not explained by the causal mechanism is considered to be noise, often framed into the formalism as an exogenous variable (Zanga et al., 2022).
In general, identifying one of the elements of the MEC requires interventions, which in this context have been formalized by Pearl (Pearl, 2009). Formally, interventions correspond to modifications of the base structural causal model by altering or replacing the causal mechanisms (i.e., the incoming functions) for one or more chosen variables. Each such intervention defines a new “environment” whose data are still drawn i.i.d. from the intervened model .
The concept of exchangeability was first used in the field of causal discovery by Guo et al. (Guo et al., 2024). The main motivation arose from the fact that, by considering the data to be exchangeable, it would allow more complex dependencies between the data, which could then be explored, even in the bivariate case.
In this paper, we argue that the advantage of considering the data to be exchangeable lies far beyond this original motivation. More specifically, it stems from the fact that it is a fairer real-world representation. In order to do so, we begin by laying out the problem and the work done so far in this field in Section 2. Then, in Section 3, we explore what are the exact implications of assuming exchangeability instead of i.i.d., while giving the relevant arguments why. Afterwards, we dive into our synthetic dataset, how it was generated, and how it compares to others in Section 4. Lastly, in Section 5, we introduce a neural network trained on this dataset, which serves both as a valid causal discovery method and a verification of the assumptions made throughout this work.
2 Problem Setting and Related Work
2.1 Exchangeability
Definition: An exchangeable sequence of random variables is a finite or infinite sequence $X_{1},X_{2},X_{3},\ldots$, such that for any finite permutation $\pi$ of the position indices $\{1,\ldots,N\}$, the joint distribution of the permuted sequences is the same as that of the original: An exchangeable sequence thus verifies:
$P(X_{\pi(1)},\ldots,X_{\pi(N)})=P(X_{1},\ldots,X_{N}).$ (1)
Additionally, a sequence $X_{1},\ldots,X_{N}$ is partially exchangeable if there exists a partition of the indices such that permutations within each partition are exchangeable.
In essence, exchangeability is a notion of symmetry (Guo et al., 2024); informally, it states that the order in which the variables appear does not matter. Naturally, all independent and identically distributed (i.i.d.) sequences are exchangeable since $P(X_{1},\ldots,X_{N})=\prod_{i=1}^{N}P(X_{i})$. However, the opposite is not necessarily true, since exchangeability doesn’t imply either independence or an identical distribution of the variables. To better understand why, consider the Pólya urn setup: a sequence of random draws of black or white balls from an opaque urn with replacement, whose proportion of black and white balls is unknown. In this scenario, each ball updates our belief over the probability distribution inside the urn. For example, if our first three draws are white balls, then it is more likely that our next draw is white and vice versa. Therefore, it breaks the independence property of i.i.d. data. However, this sequence is still exchangeable, since any finite sequence of draws from the urn provides exactly the same information regardless of the order, and therefore the different draws are indeed exchangeable. Additionally, it is also important to understand what it means for a sequence not to be exchangeable: it implies that the order of the variables matters. In other words, the observed sequence represents a time series.
The most important result regarding exchangeability is given by the famous de Finetti Theorem, stated next:
De Finetti Theorem (De Finetti, 1931):
Let $(X_{n})_{n\in\mathbb{N}}$ be an infinite sequence of binary random variables. The sequence is exchangeable if and only if there exists a random variable $\theta\in\Theta$, with probability measure $\mu$, such that $X_{1},X_{2},\dots$ are conditionally i.i.d. given $\theta$, that is, for any given any sequence $(x_{1},\dots,x_{N})\in\{0,1\}^{N}$,
$P(X_{1},\ldots,X_{n})=\int_{\Theta}\prod_{i=1}^{n}P(X_{i}\mid\theta)\,d\mu(\theta).$ (2)
In Equation (2), $d\mu(\theta)$ can be replaced by $p(\theta)d\theta$, if $\mu$ is absolutely continuous with respect to the Lebesgue measure on $\Theta$ (which is the only case of interest in this paper, and thus assumed in the sequel). The theorem has been extended in several ways, including to non-binary variables.
This theorem states that any sequence of exchangeable variables can be seen as a mixture of i.i.d. sequences. Each i.i.d. sequence is given by conditioning the data on the parameter $\theta$ and the mixture’s proportions are consequently defined by $p(\theta)$. This result agrees with the intuition from the Pólya urn examples. The case of i.i.d. data can be seen as the special scenario when $p(\theta)=\delta(\theta_{0})$. In this case, the expression simply becomes $p(X_{1},\ldots,X_{n})=\prod_{i=1}^{n}p(X_{i}\mid\theta_{0})$, which is obviously i.i.d..
Lastly, it is important to mention that it is the exact existence of this parameter $\theta$ that allows the expression of learned information about the system. For example, in the original Pólya urn, one can reframe each draw’s dependence on the previous results by stating that they are informative over this parameter and its probability distribution. Naturally though, the solution for it is not unique under an infinite sequence of exchangeable data. Nevertheless, it is possible to infer valid values for both $p(x_{i}\mid\theta)$ and $p(\theta)$ from an infinite exchangeable sequence. This has been seen by Bernardo and Smith for a sequence of a Bernoulli variable (Bernardo and Smith, 2009).
In Bayesian Theory, this $\theta$ is interpreted as a latent variable. Note that $p(\theta)$, which appears naturally through the theory of exchangeability, is exactly equivalent to our prior belief on $\theta$ in Bayesian theory. Therefore, it is not by mathematical need or lack of knowledge that Bayesian inference requires a prior on $\theta$. In fact, it stems from $\theta$ being a real latent variable which indeed follows a given probability distribution (Fortini and Petrone, 2025). It demonstrates that it is the belief that the observations are exchangeable, not any metaphysical belief about the true model, that underpins the use of Bayesian modeling involving i.i.d. observations conditioned on some unknown latent variable.
2.2 Causal Discovery
Causal discovery is the problem of finding the graph that represents the causal relationships between a collection of variables of interest (Pearl, 2009). Usually, this graph is considered to be directed and acyclic. The absence of cycles implies it cannot capture the nature of systems with feedback. This assumption is taken in order to simplify the problem, however, in some cases, causal discovery methods may still work even if the acyclicity assumption does not hold (Glymour et al., 2019). Alternatively, some methods have also been proposed that relax the assumption of acyclicity (Richardson, 2013; Lacerda et al., 2012). Additionally, both the faithfulness assumption, which means that every conditional independence property in the joint probability distribution corresponds to a separation property in the graph, and the Markov property, which states that any variable should be independent of all other variables in the graph when conditioned on its parents, should hold (Pearl, 2009).
To answer counterfactual questions (”What would have happened if…?”), one requires a more complete description of the causal system than just the information contained in a causal graph. This description can be expressed in its most general form as a structural causal model (SCM). If we restrict an SCM to a specific set of parametrized functions $f(\cdot,\cdot;\eta)$, representing the causal relationships, we obtain a functional causal model (FCM), which is defined by its DAG and the functions relating its variables,
$X_{i}=f(PA(X_{i}),\epsilon_{i};\eta_{i}),$ (3)
where $PA(X_{i})$ is the set of parents of $X_{i}$ in the DAG and each $\epsilon_{i}$ is an exogenous variable (Zanga et al., 2022).
As mentioned in Section 2, in the bivariate scenario, it is impossible to distinguish between the two possible causal directions without interventions or additional assumptions. Therefore, different methods are based on different assumptions and argue why these should only hold in the true causal direction. Some try to exploit the independence between different variables, for example, independence between noise and cause (Peters et al., 2010) or independence between noise and function mechanism (Janzing et al., 2012). Another family of assumptions is based on Occam’s razor, restricting the set of accepted causal mechanisms and choosing the causal direction that has a better fit. For example, Blöbaum et al. restrict the functions to be linear, Goudet et al. opt for a neural network with one hidden layer, while Dhir et al. use Bayesian model selection with priors on the functions to be more flexible (Blöbaum et al., 2018; Goudet et al., 2018; Dhir et al., 2023). Lastly, some methods are based on analysing the complexity of the model in each direction, assuming the factorized model in the true direction has a lower Kolmogorov complexity; however, since Kolmogorov complexity is not computable, it is necessary to resort to different proxies of this measure (Marx and Vreeken, 2017; Tagasovska et al., 2020).
Regrettably, the causal discovery research area is critically lacking in extensive real‐world datasets [^3] Here, “extensive” denotes the presence of many distinct examples. Although numerous multivariate datasets exist, each comprises only a single causal graph, which precludes the use of evaluation metrics such as accuracy., with the Tübingen dataset remaining the only one widely accepted in the literature (Mooij et al., 2016). It contains a total of 108 causes effect pairs with known ground truth based on expert knowledge. These pairs were gathered from 37 different domains; some examples are the altitude and temperature of cities, the horsepower and fuel consumption of cars, and the age and height of different people. Given the lack of good real-world benchmarks, it is common to also test new methods on synthetic datasets. For the bivariate problem in particular, there exist 4 commonly used datasets (CE-Cha, CE-Net, CE-Gauss, and CE-Multi) Guyon et al. (2019). These datasets (with the exception of CE-Cha) were designed having certain assumptions in mind (such as additive or multiplicative noise) and thus are usually used to test whether certain methods can detect the true causal direction under those specific hypotheses and not to assess the more general performance of the methods.
Lastly, causal representation learning (CRL) is an emerging field that extends traditional causal discovery to high-dimensional data by aiming to recover latent causal factors from raw observations (Schölkopf et al., 2021). It assumes the observations are generated by a set of low-dimensional latent variables that follow an SCM and seeks to learn neural encoders and decoders that map between the observations and these abstract causal variables. By embedding a causal structure into a deep learning architecture, CRL enables the model to learn disentangled and modular representations of the underlying data-generating process. Such representations are often interpretable and robust, as independent causal factors tend to remain invariant under interventions or distribution shifts, allowing model components to be reused or fine-tuned for new tasks.
2.3 Previous Work Intersecting Causality and Exchangeability
Currently, the concepts of causality and exchangeability overlap in two principal domains: the well-established field of causal inference, which fundamentally relies on exchangeability, and the recent work that generalizes de Finetti’s theorem to the causal case and introduces a novel algorithm for causal discovery (Guo et al., 2024).
In the experimental setting, often the goal is to perform cause-effect identification. This means inferring whether a variable has a causal effect on another, while knowing the reverse is not possible (for example, if one variable corresponds to an event that precedes the other). This is often assessed using randomized trials, which can be viewed through the potential outcomes framework (Rubin, 2005). Suppose one wants to assess if a certain treatment ($T$) is effective against a specific disease (outcome $Y$). To do so, a common approach is to gather a group of patients, of which half takes the treatment ($T=1$) and another half does not ($T=0$); the latter is usually called the control group. From this data, the goal is to try to estimate whether the treatment was effective or not. This is given by the average treatment effect (ATE) defined as (Saarela et al., 2023)
$\text{ATE}=\mathbb{E}_{D}[Y(1)-Y(0)]=\mathbb{E}_{D}[Y(1)]-\mathbb{E}_{D}[Y(0)].$ (4)
where $\mathbb{E}_{D}$ is the expected value of each variable over the entire distribution, and $Y(1)$ and $Y(0)$ are two random variables representing the outcome for $T=1$ and $T=0$ respectively.
The main challenge lies in computing this expression since, for each patient, it is either observed $Y(1)$ or $Y(0)$, never both. Additionally, it obviously makes no sense to consider each patient as a different i.i.d. variable of the same process, given the relevant differences between patients, such as age and health condition.
In order for ATE to be computable, exchangeability between the different patients must be assumed (Rosenbaum and Rubin, 1983). To see why, all that is needed is to manipulate the original exchangeability equality (Equation 2.1). Suppose the indices up to n belong to the control group, and the goal is to generalize the results for the treatment group in the control group. This can be done because
$P(Y_{1}(1),\ldots,Y_{n}(1),Y_{n+1}(1),\ldots,Y_{2n}(1))=P(Y_{n+1}(1),\ldots,Y_{2n}(1),Y_{1}(1),\ldots,Y_{n}(1)),$ (5)
[…]
$\iff\mathbb{E}_{\{(t,y)\in D:t=1\}}[Y(1)]=\mathbb{E}_{\{(t,y)\in D:t=0\}}[Y(1)].$ (7)
Therefore, the fact the data is exchangeable implies that $\mathbb{E}_{\{(t,y)\in D:t=1\}}[Y(1)]$ can replace $\mathbb{E}_{D}[Y(1)]$ (and similarly for the control group) (Hernán and Robins, 2020), which allows computing the ATE. More intuitively, the data being exchangeable means that it is fair to generalize the results obtained in one group to the entire population. This is valid since the patients in both groups were drawn from the same latent distribution (Höfler, 2005). In a more basic level, this also justifies the need for the patients to be picked at random, so that their choice is not informative over the underlying latent variables distribution.
Additionally, the converse is also true: if the data is not exchangeable, then it is impossible to compute the ATE, since it is impossible to generalize to the entire population. If this were the case, it would mean there was some difference in the latent variable distribution between both groups (e.g. one group could be younger than the other). This would invalidate extrapolating the results obtained from one group to the other, as this could lead to misleading results. Nevertheless, if one is aware of such imbalances, it is possible to adjust for them by considering them as observed variables (if they are available), and performing the exact same reasoning while conditioning on these observed variables (Lee and Lee, 2022).
Interestingly, within the potential outcomes framework, exchangeability implies that correlation is equivalent to causation (Rosenbaum and Rubin, 1983; Hernán and Robins, 2020). This is because under the exchangeability hypothesis, all latent variables (which could generate confounding) are controlled for by both groups following the same prior.
The connection between exchangeability and causal discovery was made only recently: in work that remains the only joining both fields, Guo et al. adapt de Finetti’s Theorem (see Section 2) to the setting of causal discovery.
Causal de Finetti Theorem (Guo et al., 2024):
Let $\{(X_{n},Y_{n})\}_{n\in\mathbb{N}}$ be an infinite sequence of binary random variable pairs. The sequence satisfies the following properties: 1) It is infinitely exchangeable; 2) $\forall n\in\mathbb{N}:Y_{[n]}\perp\!\!\!\perp X_{n+1}\mid X_{[n]}$, where $n]=\{1,\ldots,n\}$, if and only if there exist two random variables $\theta\in[0,1]$ and $\psi\in[0,1]$, with probability measures $\mu$ and $\nu$, respectively, such that the joint probability can be represented as:
$P(X_{1},Y_{1},\dots,X_{N},Y_{N})=\int_{\Phi}\int_{\Theta}\prod_{n=1}^{N}p(Y_{n}\mid X_{n},\psi)p(X_{n}\mid\theta)\,d\mu(\theta)\,d\nu(\psi).$ (8)
In (8), $d\mu(\theta)$ and $d\nu(\psi)$ can be substituted by $p(\theta)d\theta$ and $p(\psi)d\psi$, respectively, since $\mu$ and $\nu$ will always refer to absolutely continuous densities w.r.t. Lebesgue measure on $\Theta$ and $\Psi$.
Additionally, Guo et al. also proposed an algorithm for causal discovery in the bivariate setting based on the asymmetry between both causal directions present in the second assumption of the Causal de Finetti Theorem. The idea is that, while in the causal direction, $Y_{[n]}\perp\!\!\!\perp X_{n+1}\mid X_{[n]}$ holds, the same is not necessarily true in the opposite direction. To test whether $\forall n\in\mathbb{N}:Y_{[n]}\perp\!\!\!\perp X_{n+1}\mid X_{[n]}$, one would need to have the joint probability distribution over $P(Y_{[n]},X_{n+1},X_{[n]})$. However, the asymmetry expressed above is impossible to test in an exchangeable process since only one realization of each variable is obtained in reality. To go around this issue, the authors constructed datasets that approximate de Finetti’s Theorem in viewing an exchangeable process as a mixture of i.i.d. sequences. To do so, they assume the data comes from N different environments, where each environment, $e$, is defined by its latent parameters, which are drawn according to the prior distribution: $(\theta^{e},\psi^{e})\sim(p(\theta),p(\psi))$. In a sense, this is similar to the earlier concept of interventions, because all environments share the same underlying causal graph over the observed variables $X$ and $Y$, while the specific causal mechanism is defined by $\theta$ and $\psi$. Nevertheless, from a more practical viewpoint, these interventions are much softer between each other (very similar causal relationships between interventions), and many more in number (the author’s algorithm considers 100).
3 Exchangeability in Causal Discovery
This paper aims to recast i.i.d. causal discovery in terms of exchangeable data. We begin by examining the potential consequences of this reformulation and then offer a series of arguments to justify its adoption.
3.1 Implications
Since exchangeable sequences may not be i.i.d., it is important to consider what properties and differences exchangeable sequences hold in general, which are absent in i.i.d. sequences.
The first relevant difference is that, in exchangeable sequences, the different samples may not be independent from each other. Guo et al. proposed a causal discovery algorithm by noticing that a certain conditional independence relationship between different samples only holds in one causal direction (Guo et al., 2024), as explained in Section 2.3. New methods for causal discovery could be developed by identifying and mathematically formalizing asymmetries in data that are induced by the direction of the underlying causal relationship. Additionally, it is also relevant to point out that all causal discovery methods that are valid for the i.i.d. case should also be valid for exchangeable data. This stems from these methods relying upon an even more restrictive assumption, which holds in the causal direction for i.i.d. data; therefore, it is naturally generalizable to exchangeable data. On the other hand, the opposite is not necessarily true. Causal discovery methods developed for exchangeable data might take advantage of asymmetries that disappear when we focus on i.i.d. data (see Section 2.3). Therefore, considering the data to be exchangeable instead of i.i.d. opens the door to new methods that may take advantage of new patterns in the causal relationships, while ensuring the verified methods for i.i.d. data keep their validity.
From a different angle, the de Finetti theorem (see Section 2.3) states that any exchangeable sequence can be written as a continuous mixture of i.i.d. sequences. This mixture is controlled by unobserved latent variables. At first glance, these latent variables might seem to be exactly the same as the exogenous variables in the SCM and FCM models, which are usually interpreted as noise; in fact, mathematically speaking, this is the case. However, noise is traditionally used with a very different connotation from the notion of latent variables. For one, noise tends to be thought of as having a small effect in the data, whilst the existence of a latent variable is generally assumed to be as significant as the observed ones. Secondly, noise is also generally believed to affect the causal function in simple ways (one often considers noise to be additive or multiplicative), whilst a latent variable is considered to introduce significant complexity in this relationship. Therefore, even though there is no strict difference between naming the exogenous variables as noise or latent variables, there are many design disparities that may stem from this difference. Consequently, considering the data to be exchangeable would imply the existence of latent variables, which could motivate causal discovery methods to ponder their existence explicitly and thus take into account the variety of complex effects they might have on the observed joint distribution.
Lastly, it is interesting to note that exchangeability between samples can also be linked with the idea of interventions (see Section 1). One can think of an exchangeable sequence as corresponding to each sample being obtained from a different environment (different intervention). In this scenario, each environment would be defined by the latent variable, which in turn changes the causal relationship between the observed variables.
3.2 Arguments
3.2.1 Conceptual Argument
The conceptual argument in favor of framing i.i.d. causal discovery in exchangeable data starts by noticing the clear similarities between causal inference and causal discovery. Crucially, they share the same setup and the same ultimate goal. The difference is that in causal inference, it is possible to intervene directly in the system, while causal discovery relies solely on observational data. Since they operate under the same formal assumptions and objectives, it makes sense to require the same foundational properties of the data in both settings. In experimental causal inference, one of the key foundational assumptions is exchangeability, which allows for controlling for external factors (see Section 2.3). It has been widely accepted and thoroughly validated in the scientific community as the criterion that enables extrapolating causal conclusions drawn from experimental data (Saarela et al., 2023).
While experimental causal inference has long been a routine, rigorously validated practice, observational causal discovery remains primarily an exploratory research field. Consequently, observational causal discovery should look towards causal inference in order to extrapolate verified and meaningful assumptions on the data. More specifically, it should resemble causal inference in considering exchangeability to be the key statistical property of observational data.
This perspective contrasts sharply with the traditional assumption that the observational data are independent and identically distributed (i.i.d.), since this would require that the system is completely isolated. In turn, the introduction of exchangeability opens the possibility of considering the existence of external unobserved variables. This precise existence is supported by causal inference posing exchangeability as one of its foundational assumptions. Consequently, a failure to recognize such latent variables will necessarily result in the development of extremely narrow-sighted causal discovery algorithms.
3.2.2 Empirical Argument
The empirical argument for reframing i.i.d. causal discovery using exchangeability draws on recurring patterns in recent causal-discovery research that become clearer under this lens. First, we re-examine the Tübingen dataset and show that most of its examples are better characterized as exchangeable rather than strictly i.i.d. Then, we survey state-of-the-art methods to identify those that implicitly exploit exchangeability, typically by modulating the latent variable at the heart of de Finetti’s theorem.
As mentioned in Section 2.2, the Tübingen dataset collection is the only extensive real-world i.i.d benchmark present in the literature and widely used to assess bivariate causal discovery methods. Consequently, we decided to analyze it closely, in order to understand whether the examples it includes are strictly i.i.d. or whether their samples should be considered exchangeable. From the de Finetti theorem (Section 2), exchangeability implies the existence of latent variables, whilst i.i.d. data only occurs when the system is isolated. Therefore, we checked if there existed any plausible unobserved variable in each example of the dataset (the full classification can be seen in Appendix A). After doing so, we concluded that the vast majority of the datasets in the collection ($81.5\%(74.4\%)$ [^5] The second result corresponds to the weighted average, according to the weights proposed by the authors of the dataset.) should be actually considered as exchangeable, while there was a small percentage containing timeseries ($10.2\%(15.4\%)$) and another that can considered i.i.d. ($8.3\%(10.3\%)$).
Several exchangeable examples can be seen in datasets relating geographical features (either latitude, longitude, or altitude) with meteorological quantities (temperature and precipitation). In this scenario, the hidden geographical features are clearly latent variables. A similar situation happens in the datasets connecting the compressive strength of cement with each of its mixture components, or the fuel economy of cars with each of its many features. Additionally, there are many other examples that have very reasonable latent variables. The sets containing the associations between different features of living beings (humans and oysters in specific) with their age clearly involve latent variables in the genetics of each individual.
On the other hand, there are two sets of examples that tried to eradicate the influence of latent variables. One relates different physical attributes of a rolling ball (such its height), with the initial and final velocity, while the other contains both measurements of the emitted and perceived light by a block of LEDs and a photo receptor.
Finally, some datasets fall in the category of timeseries, such as those relating inside and outside temperature or those that associate temperature with the day of the year. As we know, timeseries are not exchangeable, so these could undermine the conclusions we were trying to reach. However, since causal discovery with i.i.d. data is a slightly separate field of research from in timeseries, we believe these examples are present in this dataset by mistake. In principle, any method that works well on i.i.d. or generally exchangeable data should also perform well on timeseries. In practice, however, explicitly recognizing the sequential nature of timeseries lets us design algorithms that exploit far richer temporal dependencies—and ultimately uncover more nuanced causal relationships. The results of our analysis of the Tübingen dataset are shown in Figure 1.
Figure 1: On the left, we can see the unweighted and weighted distribution of statistical assumptions, while on the right, we have 2 examples of timeseries present in the Tübingen dataset.
[…]
The second implication of reframing i.i.d. causal discovery as exchangeable pertains to the difference in interpreting unobserved variables as latent variables instead of noise (see Subsection 3.1). Therefore, methods that allow the unobserved variables to influence the effect in complex and strong ways, can be seen as assuming the existence of latent variables. This allows them to be interpreted in the light of exchangeability, not exactly satisfying the i.i.d. assumption in this area of causal discovery.
The cleanest example in the state of the art is CGNN (Goudet et al., 2018). This method chooses the causal graph whose generated distribution using a neural network lies closest to the true distribution. Both observed and unobserved variables are considered as inputs to the neural network. Therefore, there is complete symmetry between what is observed and unobserved (the only difference is that the marginal distribution of observed variables is obtained from the data, while that of unobserved variables is assumed according to prior knowledge). Therefore, the structure of the neural network ensures that the unobserved variables can contribute in exactly the same way to the effect as the observed causes.
Another interesting approach to exchangeability is bCQD (Tagasovska et al., 2020), which attempts to perform causal discovery by comparing the complexity of enconding the joint distribution using the two possible causal factorizations [^7] The two possible causal factorizations are: $P(X,Y)=P(Y\mid X)P(X)$ and $P(X,Y)=P(X\mid Y)P(Y)$.. However, and more relevant to the subject of this paper, it models the function as a set of quantile regressions. The underlying rationale can only be understood if we consider the existence of a latent variable, and each different quantile regression is trying to model the system for a fixed value of the latent variable. Therefore, assuming exchangeability also contributes to a clearer understanding of this method.
It is important to mention, however, that while some methods are much better understood by extrapolating to outside the i.i.d. domain, this is not generally the case. There are also a variety of methods whose main assumptions are placed in the nature of the causal function, such as SLOPE (Marx and Vreeken, 2017) and IGCI (Mian et al., 2023). In other methods, the noise nature of the unobserved variables (weak impact and low complexity) are a key part of the algorithm (see RECI (Blöbaum et al., 2018) or LiNGAM (Shimizu et al., 2006)). Nonetheless, this is to be expected given how the causal discovery problem is formulated for i.i.d. data. We believe that the emergence of methods based on exchangeability within this domain, achieving performance comparable to other state-of-the-art algorithms [^8] See Guyon et al. and Tagasovska et al. for a comparative analysis of different methods(Guyon et al., 2019; Tagasovska et al., 2020)., serves as further evidence supporting our central argument. Moreover, we are hopeful that extending i.i.d. causal discovery to the exchangeability framework will open up opportunities for many additional methods to be developed.
4 Synthetic Dataset
4.1 Motivation and Objectives
The creation of this synthetic dataset of cause–effect pairs fulfills two objectives simultaneously:
  • Provide an additional, valid dataset for testing new methods: as noted in Subsection 2.2, the Tübingen dataset remains the sole real-world benchmark for causal discovery. Existing synthetic datasets, in contrast, are used only to evaluate method validity under particular assumptions. The goal of our dataset is to offer a properly constructed synthetic resource for method analysis. We propose using it complementarily alongside the Tübingen dataset: since it is synthetic, it allows for far greater precision (examples can be generated as needed) and excludes certain questionable instances (see Subsection 3.2.2 regarding time-series examples and independence considerations).
  • Showcasing the central role of exchangeability in causal discovery: this dataset is created with a focus on exchangeability. This is achieved by generating samples according to the Causal de Finetti theorem (see Subsection 2.3). With the goal of ensuring exchangeability takes central stage in our dataset, we randomize the choice of both prior distributions and causal functions across different examples, using widely accepted classes. Therefore, by doing so, we aim at showing that if this dataset is able to mimic the real-world dynamics, then exchangeable (not i.i.d.) distributions must be at the center of causal discovery.
To guarantee both objectives above, it is necessary to demonstrate that the dataset is valid. In other words, it must be shown to be representative of real-world scenarios. Because the only available real-world reference is the Tübingen dataset, this can be achieved by demonstrating that various state-of-the-art methods produce very similar results on both the Tübingen dataset and the proposed synthetic dataset.
4.2 Algorithm
The algorithm behind the synthetic dataset generation is based on the Causal de Finetti theorem (see Subsection 2.3). According to Equation 8, in order to construct a bivariate dataset satisfying exchangeability, one needs to model the distribution of both latent variables ($p(\theta),p(\psi)$) and the causal functions which are represented by the two conditional probabilities ($p(y_{n}\mid x_{n},\psi),p(x_{n}\mid\theta)$). Given that $x_{n}$ only depends on $\theta$, and the generative nature of our objective, the conditional function relating the two can be fully expressed by taking $\theta$ to represent the true distribution of x, and the conditioning only adding noise to the measurement. On the other hand, $p(y_{n}\mid x_{n},\psi)$ is decomposed into the true causal function $y_{true}=f(x_{n},\psi)$ and a Gaussian noisy measurement $y_{n}=\mathcal{N}(y_{true},\sigma^{2})$ [^9] Note that in both conditional probabilities, we can force the noise power to be zero ($\sigma^{2}=0$) which would yield a noiseless version of the same exchangeable system., with variance $\sigma^{2}$. Therefore, a causal exchangeable dataset generation process can be implemented using Algorithm 1.
Algorithm 1 Synthetic Dataset Generation Algorithm
Input: $p(\theta),p(\psi),f(),N_{S},\sigma_{x},\sigma_{y}$
   Output: $N_{S}$ causally related exchangeable pairs $\{(x_{i},y_{i})\}$
[…]
2:Perform scaling on both $\theta$ and $\psi$.
[…]
6:Perform min-max scaling on both $X$ and $Y$.
Additionally, to generate many examples, it suffices to wrap Algorithm 1 within an outer loop. In the current implementation, the following considerations are made over the general algorithm and the required hyperparameters and functions (further details about specific design choices made for each causal function can be found in Appendix B.1.)
  • The priors $p(\theta)$ and $p(\psi)$ are chosen at random from 3 different distributions: Uniform, Normal, and Rayleigh. These distributions were chosen because: 1) each distribution captures common patterns observed in empirical data (e.g., a roughly uniform spread of ages in a population or the annual temperature profile approximating a Gaussian). 2) Uniform, Normal, and Rayleigh distributions form foundational building blocks across fields—from entropy-based analyses in information theory to error modeling in signal processing—and underpin key results such as the Central Limit Theorem. Their ubiquity ensures that conclusions drawn under these priors can be readily compared and contextualized with existing work. 3) Once each distribution is linearly scaled to the interval $[0,1]$, no additional shape or scale parameters remain to be tuned, thereby simplifying posterior sampling without introducing extra hyperparameter dependencies.
  • The function $f$ should model common causal relationships in the real-world. Therefore, so far, eight different functions are implemented: 1) linear; 2) piecewise linear; 3) exponential; 4) logarithm; 5) inversely proportional; 6) Brownian-like motion  [^10] Note that this function, despite requiring knowledge of previous samples to compute the latter, does not constitute a timeseries, given there exists no underlying time variable, modeling both $x$ and $y$. Alternatively, it only aims at generating a differentiable function in its most general form.; 7) polynomial; 8) power law. It is also important to point out that all these functions are designed to be strictly increasing in the latent variable ($\psi$). Although this is not directly enforced from the theory of exchangeability, it constitutes a design choice based on theoretical and practical reasons. On the one hand, monotonic SCMs have been proposed as a reasonable simplification of the problem of causal discovery (Izadi and Ester, 2024). On the other hand, many real-world examples in the Tübingen collection (see Appendix A) reinforce this belief. Examples of this are the relationship between age and height, or monthly rent and size in square meters [^11] In the first example, the latent variable can be considered to be the genes; in this case, it is possible that someone with ”taller” genes is both taller than other people at age 10, 40, 70, and so on. In the second example, considering a latent variable such as location, it is also reasonable to assume that independently of the size in square meters of an apartment, the better its location, the more expensive the price..
  • The variances $\sigma_{x}^{2}$ and $\sigma_{y}^{2}$ are small because they are designed to represent noise present in the system.
  • The number of points in each example is sampled from a distribution that aims to mimic the distribution present across the Tübingen dataset. This is obtained by fitting a Gaussian mixture with 3 components to the data.
4.3 Constructing the Full Dataset
When constructing the full dataset, there exists an additional aspect to consider. In Algorithm 1, three components have been left as design options: the choice of function $f$ and the a priori distributions of the latent variables $p(\theta)$ and $p(\psi)$. Although any fixed choice of these elements yields an exchangeable synthetic dataset, the appropriate frequency of each specification within the global synthetic dataset should be adjusted to mirror the global structure of real data.
To address this issue, nine different established causal discovery methods were implemented: ANM (Peters et al., 2010), bQCD (Tagasovska et al., 2020), CGNN (Goudet et al., 2018), EMD (Chen et al., 2014), IGCI (Mian et al., 2023), LiNGAM (Shimizu et al., 2006), PNL (Zhang and Hyvärinen, 2010), RECI (Blöbaum et al., 2018), and SLOPE (Marx and Vreeken, 2017) [^12] The implementations of RECI and IGCI were taken verbatim from Kalainathan and Goudet (Kalainathan et al., 2020). CGNN and ANM are based on the same repository, with minor adjustments to their hyperparameters (for ANM, we replaced its original independence test with the Hoeffding’s D test to improve computational efficiency). The bQCD and EMD methods were sourced from Tagasovska et al., and SLOPE was implemented as described in its original publication (Tagasovska et al., 2020; Marx and Vreeken, 2017). Finally, the LiNGAM and PNL algorithms were obtained from the GitHub repository at https://github.com/ssamot/causality/tree/master.. Afterwards, their performance in two different metrics, AUROC (area under the ROC curve) and accuracy [^13] These two metrics are widely present in the literature, hence their use in this work., was recorded for each combination of the three existing design choices. In Table 1, partial results obtained by these nine methods are reported. In the absence of any prior bias, the generated dataset was uniformly sampled from across the different combinations of design choices mentioned previously. Consequently, the results displayed in the table can be seen as the average obtained for each function across the different combinations of the choice for the latent variables’ distribution.
Method All Brownian-like Exponential Inv. Proportional Linear Logarithmic Piecewise Polynomial
ANM 0.529 (0.512) 0.248 (0.296) 0.768 (0.703) 0.256 (0.251) 0.538 (0.514) 0.785 (0.771) 0.232 (0.277) 0.194 (0.256)
bQCD 0.420 (0.421) 0.500 (0.494) 0.463 (0.453) 0.570 (0.531) 0.354 (0.365) 0.246 (0.286) 0.468 (0.456) 0.573 (0.511)
CGNN 0.682 (0.653) 0.635 (0.624) 0.777 (0.722) 0.727 (0.673) 0.583 (0.565) 0.854 (0.810) 0.605 (0.591) 0.559 (0.562)
EMD 0.737 (0.718) 0.667 (0.655) 0.964 (0.906) 0.811 (0.785) 0.551 (0.574) 0.969 (0.938) 0.559 (0.575) 0.584 (0.575)
IGCI 0.758 (0.726) 0.704 (0.675) 0.952 (0.896) 0.836 (0.789) 0.594 (0.599) 0.986 (0.939) 0.583 (0.593) 0.588 (0.592)
LiNGAM 0.508 (0.515) 0.466 (0.462) 0.644 (0.612) 0.567 (0.574) 0.315 (0.365) 0.756 (0.701) 0.363 (0.419) 0.403 (0.432)
PNL 0.478 (0.491) 0.533 (0.531) 0.452 (0.485) 0.487 (0.496) 0.473 (0.487) 0.449 (0.477) 0.495 (0.492) 0.487 (0.480)
RECI 0.742 (0.732) 0.678 (0.686) 0.867 (0.876) 0.851 (0.821) 0.611 (0.630) 0.975 (0.941) 0.615 (0.619) 0.614 (0.596)
SLOPE 0.757 (0.729) 0.683 (0.665) 0.967 (0.922) 0.828 (0.803) 0.610 (0.631) 0.953 (0.906) 0.624 (0.642) 0.609 (0.600)
Table 1: Marginalized AUROC (accuracy) achieved by different state-of-the-art methods in a uniformly sampled exchangeable synthetic dataset.
Since the algorithms should perform similarly in this dataset as with the Tübingen dataset, in order for it to fulfill its objectives (see Subsection 4.1), the frequency of each combination of parameter choices should be determined such that the results obtained in the two different datasets are as similar as possible. Alternatively, instead of attributing different frequencies to each combination, each one was attributed a different weight (assuming a uniformly sampled distribution across all combinations, i.e., all combinations have the same number of examples). This provides extra flexibility, since different weights can be chosen with different objectives in mind, without the need to generate different examples. Specifically, two different weight combinations are provided in our implementation: one aimed at minimizing the difference in performance with respect to AUROC, the other with respect to accuracy. Both were computed by framing the problem using least squares optimization. In other words, it consists of finding the weight combination that minimizes the sum of the squared difference between each method’s performance in both datasets. Additionally, an $\ell_{2}$ regularizer was also used in order to improve performance for unseen methods [^14] The $\ell_{2}$ regularizer and its weight ($1$) were chosen to optimize the results of SynthNN (see Section 5). In other words, the performance of SynthNN was used as a proxy for the representativeness of the synthetic dataset.. This can be translated into the following mathematical expression:
$\min_{w}$ $\|A\,w-b\|_{2}^{2}\;+\;\,\|w\|_{2},$ (9)
subject to $w\geq 0,$
$\mathbf{1}^{\top}w=1,$
where $A\in\mathbb{R}^{m\times d}$ is a matrix where element $A_{i,j}$ corresponds to the performance of method $i$ in the generated dataset according to the design choices $j$; $b\in\mathbb{R}^{m}$ is a vector containing the performance of the assessed methods in the Tübingen dataset; $w\in\mathbb{R}^{d}$ is a vector containing the corresponding weight of each parameter combination.
Lastly, AUROC is a ranked choice metric, which means it analyses the order of confidence between different guesses. Therefore, the equation above does not exactly optimize the average error between datasets in the AUROC metric. This is because the weighted average of the AUROC for each design choice combination isn’t the AUROC of the weighted dataset. Nevertheless, optimizing the AUROC exactly is a combinatorial problem, much harder to solve, making the expression above a useful and simple approximation.
4.4 Analysis
In order for the two objectives laid out initially to be fulfilled, the algorithms should perform similarly in the synthetically generated dataset presented in this paper as in the Tübingen dataset.
Firstly, it is important to understand which shapes the different generated examples can take. The different causal functions presented in Subsection 4.2 were designed to be as general and representative of the real world as possible. Figure 2 shows how synthetic examples can match those in the Tübingen dataset by fixing its random operations. Therefore, it seems reasonable to conclude that the data‑generation mechanism is sufficiently expressive to capture the variety of relationships found in the Tübingen dataset. Nevertheless, despite the ability of the algorithm to generate examples mimicking those in the Tübingen dataset, the inherent randomness present in the process ensures the actual examples are, in practice, much different (a random selection of these examples can be seen in Appendix B.2). It is also important to note that despite this apparent ability to represent the variety of causal functions present in the real-world, the Tübingen dataset has examples with discrete, categorical, or multi-dimensional variables, which are not present in the current implementation of the synthetic dataset.
Figure 2: Three normalized Tübingen pairs plotted alongside hyperparameter-tuned samples from the synthetic dataset proposed in this paper.
Even though the proposed data generation process seems to have the ability (through fine-tuning) to capture real-world examples, this is not enough to guarantee the quality of the dataset for two different reasons. For one, even though the generation process clearly passes the human visual test in terms of expressibility, this does not ensure it provides a proper representation of the causal element inherent to the real-world data. Secondly, even if the algorithm was demonstrably general enough to represent the causal dynamics of all possible known real-world pairs via fine-tuning, this would not necessarily extrapolate accordingly to randomly generated examples. In other words, the noticed causal element present in the fine-tuned examples might be a consequence of the fine-tuning itself, rather than the merits of the data generation algorithm.
A more comprehensive evaluation of the generated dataset was therefore undertaken. Nine established causal‐discovery algorithms were tested on both the Tübingen and our dataset, with performance measured by AUROC and accuracy. To facilitate a broader comparison, these methods were also applied to five additional synthetic datasets (including a noisy variant of our own). The full set of results is presented in Table 2.
Method Metric Ours Tübingen CE-datasets
Original Noisy Gauss Net Cha Multi
ANM AUROC 46.4 (44.0) 45.6 (45.1) 44.7 12.1 21.0 35.2 74.0
accuracy 41.0 40.2 39.7 19.3 29.3 37.7 61.3
CGNN AUROC 68.9 (67.1) 66.6 (64.9) 66.7 70.6 67.1 62.5 93.6
accuracy 61.8 61.0 62.5 64.7 62.0 61.3 85.0
EMD AUROC 78.4 (70.3) 80.4 (72.1) 69.3 57.3 71.3 59.3 98.1
accuracy 63.4 66.8 61.7 54.7 65.0 56.3 91.0
IGCI AUROC 80.8 (75.1) 80.4 (73.2) 70.7 43.0 58.9 56.9 97.8
accuracy 68.5 67.6 65.1 44.7 55.3 55.0 92.3
LiNGAM AUROC 49.6 (50.7) 50.8 (50.7) 50.0 29.9 63.4 45.0 33.9
accuracy 50.5 51.6 49.2 36.3 61.3 46.0 40.3
PNL AUROC 47.1 (47.9) 46.3 (46.0) 41.3 44.6 51.4 48.1 42.2
accuracy 48.2 49.3 44.5 45.7 49.7 48.0 45.7
bQCD AUROC 61.3 (63.5) 58.7 (61.1) 73.0 50.9 92.2 58.7 56.6
accuracy 60.2 57.1 70.0 55.3 84.0 58.3 50.3
RECI AUROC 78.6 (74.9) 78.2 (75.4) 73.9 76.5 62.8 57.7 95.4
accuracy 70.6 70.5 70.2 67.7 55.7 55.0 88.0
SLOPE AUROC 82.1 (76.4) 82.3 (76.0) 78.9 73.1 66.9 59.4 96.9
accuracy 70.6 69.8 71.5 67.3 62.3 57.0 88.7
Table 2: Performance of different causal discovery methods (in terms of AUROC and accuracy) in the paper’s dataset, the Tübingen, and five additional synthetic datasets. In the ”Ours” datasets, the extra AUROC in parentheses refers to the weighted average AUROC obtained from Equation 9
In order to compare the results shown in Table 2, the average $\ell_{1}$ and $\ell_{2}$ distances to the Tübingen dataset were computed (which can be seen in Table 3). However, since this metric was used for computing the optimal weight distribution for this paper’s dataset, it would be unfair to just present this metric, given the clear circularity in the argument: 1) The dataset quality can be assessed by how similar different methods perform when compared with the real-world reference (the Tübingen dataset) 2) There are certain design choices in the generation process that can be fine-tuned to ensure that a selection of nine different causal discovery methods have very similar performances in both datasets. 3) Since these nine causal discovery methods have very similar performances in both datasets, the data-generation algorithm (and the consequent synthetic dataset) is shown to be very good. By finetuning the dataset using the performances of some causal discovery methods, there is clear confounding in then using the same methods to assess the general similarity in performance of all causal discovery methods in the two datasets. Therefore, to control for this circularity, the average leave-one-out cross-validated $\ell_{1}$ and $\ell_{2}$ distances for this paper’s dataset were also computed. Essentially, it boils down to (one by one) first computing the weights by considering all but one method, and then analyzing the difference in performance only on the left-out method.
Norm Metric Ours (Original) Ours (Noisy) CE-Datasets
Produced CV Produced CV Gauss Net Cha Multi
$\ell_{1}$ AUROC 0.0544 (0.0297) 0.0825 (0.0613) 0.0551 (0.0322) 0.0809 (0.0627) 0.1444 0.1150 0.1104 0.2056
accuracy 0.0256 0.0519 0.0351 0.0623 0.0949 0.0878 0.0743 0.1836
$\ell_{2}$ AUROC 0.0044 (0.0018) 0.0169 (0.0115) 0.0053 (0.0021) 0.0156 (0.0107) 0.0326 0.0180 0.0146 0.0496
accuracy 0.0014 0.0083 0.0026 0.0104 0.0144 0.0098 0.0082 0.0405
Table 3: Performance of different synthetic datasets measured by comparing the average $\ell_{1}$ or $\ell_{2}$ norms of the difference between the achieved results (AUROC or accuracy) in the respective synthetic dataset and the Tübingen dataset.
Finally, as can be seen in Table 3, the presented dataset performs consistently better than the four current synthetic datasets present in the literature (CE-Gauss, CE-Net, CE-Cha and CE-Multi). The cross validated original version of the synthetic dataset surpasses all other datasets in the $\ell_{1}$ metrics (AUROC: $0.0775$ and accuracy: $0.0516$) and is only nearly outperformed by the CE-Cha dataset in the $\ell_{2}$ metrics (AUROC: $0.0120$ and accuracy: $0.0062$). However, CE-Cha was built for the cause-effect pairs challenge and contains both artificial and real-world data from the Tübingen dataset, which causes clear confounding in this analysis. Furthermore, in absolute terms, the average difference in performance between our dataset and the Tübingen is also quite small. Therefore, it is clear that the presented synthetic dataset correctly mimics the Tübingen one, and consequently, the real-world dynamics accurately. It is also clear that it does so to a higher degree than all other known synthetic datasets.
Additionally, the ”Noisy” dataset was constructed by applying significant additive and multiplicative noise to the original version, and it yields results that are marginally inferior to those of its unaltered counterpart. This suggests that there is nothing inherently causal in the presence of noise that cannot be captured appropriately by the underlying exchangeability in the dataset. This is even more interesting given the absence of any additive or multiplicative relationship between the cause and the latent variable for all implemented causal functions (see Appendix B.1).
These findings become all the more striking when we recall that exchangeability was the only design principle guiding the construction of this dataset. Even more, despite being able to achieve very similar examples to the continuous ones in the Tübingen dataset, its actual generated examples are quite different (as can be seen in Appendix B.2). This is especially the case for variables that are not continuous or one-dimensional. Consequently, the presented dataset successfully fulfills its two initial objectives (see Subsection 4.1):
  • It is clear this dataset resembles more closely the real-world than all other known synthetic datasets (see Table 3). Therefore, we hope its development will aid in better classifying, assessing, and comparing different causal discovery methods. More specifically, its complementary use may allow to compensate for some shortcomings of the Tübingen dataset, especially allowing for a more precise and general evaluation.
  • By allowing exchangeability to drive its development, we have produced the strongest synthetic dataset to date. This success lends weight to the central thesis of our paper: that causal discovery under the i.i.d. assumption should, in fact, be reframed in terms of exchangeability.
5 SynthNN
Having created the synthetic dataset explained in Section 4.1, a new causal discovery method named SynthNN is proposed that simply consists of a neural network trained on the generated data.
In each example, the data consists of a table, where there is a connection between each pair, but there is no inherent order between samples. This is given by the defining properties of exchangeable data (and consequently of i.i.d. as well). Consequently, it is not clear at first how the input data should be fed into the neural network. Therefore, four different strategies were tested:
  1. 1.
    Feeding the data as a flattened one-dimensional vector into a fully connected layer: This is equivalent to ignoring all structure existing in the data. Essentially, the relationship between each $(x_{i},y_{i})$ pairing is not enforced. This means the network has to learn (in whichever way it finds appropriate) the structure during training.
  2. 2.
    Perform feature extraction for each pair, pool all the features together, and process this feature vector; this approach is based on the PointNet architecture developed for 3D point cloud classification (Qi et al., 2017). Essentially, a shared MLP is applied individually to each pair $(x_{i},y_{i})$, which generates a higher order feature vector. Then, a global maxpooling is applied to this vector, which can be interpreted as keeping the most relevant aspects of each pair and, consequently, of the input data overall. Finally, this global feature vector is processed by another MLP.
  3. 3.
    Start by constructing a graph and then applying methods from graph neural networks (GNN) to the problem at hand. Specifically, the graph is constructed using k-nearest neighbours; then, several graph convolutional layers are applied according to the implementation by Kipf and Welling (Kipf and Welling, 2017). Afterwards, the extracted features are flattened and processed by an MLP.
  4. 4.
    Building an image from the data and processing it using a convolutional neural network (CNN). This image corresponds exactly to the images shown when plotting the generated data. Out of the four, this is the only method that was applied to the problem of classifying cause-effect pairs by Singh et al. (Singh et al., 2017). Therefore, the designed network is generally based on this earlier implementation.
At first look, the second and fourth methods have a clear advantage: they are inherently invariant in the number of pairs, which is crucial to the classification problem at hand. Alternatively, the other two would require padding so that the dimension of all examples matches. However, after implementing the four different methods, the fourth, based on image processing, clearly stood out as the most promising one. Consequently, the others were dropped and the focus from now on will be on this one. The full details of the SynthNN are in Appendix C.
The key idea behind SynthNN is that if the synthetic dataset is representative of the Tübingen collection, by training a neural network on the former, it can achieve good results in the latter. However, these displayed significant variance based on the random weight initialization performed by the neural network. Consequently, the distribution of the obtained AUROC and accuracy in the training, validation, and Tübingen sets can be seen in Figure 3.
Figure 3: Distribution of results obtained by the neural network trained in the synthetic dataset (SynthNN). First row: contains the AUROC and accuracy obtained on the training set. Second row: results on the validation set (note that this is simply a different split from the synthetic dataset). Third row: distribution of results on the Tübingen dataset.
The training and validation metrics displayed in Figure 3 show that the network is consistently learning the correct causal structure present in the synthetically generated examples. However, the performance obtained in the Tübingen data are far more varied, with AUROC having a standard deviation of $2.0\%$ and accuracy of $3.2\%$. This variation can be attributed to the weight initialization in the beginning of the training process. Nevertheless, if the training data were fully representative of the Tübingen dataset, then the random initialization would only serve to break the network symmetries and would always converge to the same result.
[…]
Furthermore, Singh et al. (Singh et al., 2017) trained a very similar neural network directly on the Tübingen dataset, having achieved an AUROC of $76.9\%$ and an accuracy of $73.3\%$. Naturally, these results represent the extent to which a neural network (with this architecture) can properly learn the causal nature present in the Tübingen data. Consequently, it should represent an upper bound for the possible performance of any similar neural network trained on synthetic data. Therefore, since the results obtained by SynthNN are not far from those obtained by training directly on the Tübingen dataset, it is reasonable to conclude that most of the remaining error should be attributed to limitations of the neural network itself.
On the other hand, it also means the synthetic dataset should represent the dynamics of the Tübingen data accurately, at least as far as the neural network is able to capture the causal relationship between data. One can think of SynthNN as offering a lower bound on how representative the synthetic dataset is of the Tübingen data and thus hopefully of the real world.
Regardless of the point of view, it seems clear that the results obtained by SynthNN provide additional evidence of how representative the developed synthetic dataset is of real-world data. As such, this method serves as further validation not only of the dataset itself but also of its underlying assumption: the central role of exchangeability in causal discovery. This is even stronger, taking into account the many unintentional additional assumptions present in the design choices of both the synthetic dataset generation algorithm and the neural network architecture.
Furthermore, SynthNN is a causal discovery method in full right, since it estimates the causal direction of real-world data, only based on first principles. This is in stark contrast to the CNN from Singh et al. (Singh et al., 2017), which is itself trained on the Tübingen data. However, unlike other methods, such as LiNGAM or bQCD, it does not have such a direct downstream application to the problem of causal representation learning. While LiNGAM and bQCD’s assumptions have concise mathematical representations and thus can be more easily fit into a machine learning pipeline, SynthNN’s key assumption lies in the representational power of the synthetic dataset, which is harder to integrate. Nevertheless, SynthNN clearly shows the validity of developing a causal discovery method by training a neural network (or any other machine learning algorithm) on a dataset that is believed to represent real-world data well.
In conclusion, we believe SynthNN shows itself not only to be a valid method for causal discovery, but also further proves the central thesis of this paper: that i.i.d. causal discovery should be reframed as exchangeable.
6 Conclusion
We have shown that the traditional i.i.d. framework for causal discovery is most naturally and more powerfully understood as exchangeable. This was first substantiated through both conceptual and practical arguments (notably, the canonical real‑world benchmark, the Tübingen cause–effect pairs, aligns better with exchangeability than strict i.i.d.). Afterwards, to make these ideas more concrete, we introduced a novel synthetic dataset that enforces only exchangeability—eschewing stronger i.i.d. constraints—and showed that its statistical properties align more closely with the Tübingen data than any prior synthetic benchmark. Complementing the dataset, we presented a neural‑network–based causal‑discovery algorithm trained solely on our exchangeable examples. This model performs similarly to other relevant and current methods on real‑world data, demonstrating the viability of an exchangeability‑only approach. Even though the broader implications of adopting exchangeability in causal‑discovery practice remain open to be fully explored, this work opens the door to novel dependencies, structures, and methodologies that more accurately reflect the complexities of real‑world data.
Appendix A Tübingen Statistical Assumptions Analysis
Table 4: Tübingen Analysis Data
Cause Effect Latent Variable Examples Exchangeable?
altitude temperature (average over 1961-1990) longitude, latitude yes
altitude precipitation (yearly value averaged over 1961-1990) longitude, latitude yes
longitude temperature (averaged over 1961-1990) altitude, latitude yes
altitude sunshine (yearly value averaged over 1961-1990) longitude, latitude yes
Oyster age (estimated using its rings) Longest shell measurement genes yes
Oyster age (estimated using its rings) Shell weight genes yes
Oyster age (estimated using its rings) Diameter genes yes
Oyster age (estimated using its rings) Height genes yes
Oyster age (estimated using its rings) Whole weight genes yes
Oyster age (estimated using its rings) Shucked weight genes yes
Oyster age (estimated using its rings) Viscera weight genes yes
Age Wage per hour education yes
displacement mpg horsepower, weight yes
horsepower mpg weight, displacement yes
weight mpg horsepower, displacement yes
horsepower acceleration weight, displacement yes
Age Dividends from stock education yes
age of child in years concentration of GAG genes yes
duration of erruption in minutes time to the next erruption in minutes weather yes
latitude temperature (averaged over 1961-1990) altitude, longitude yes
longitude precipitation (yearly value averaged over 1961-1990) altitude, latitude yes
age height genes yes
age weight genes yes
age heart rate genes yes
cement compressive strength other mixture components yes
blast furnace slag compressive strength other mixture components yes
fly ash compressive strength other mixture components yes
water compressive strength other mixture components yes
superplasticizer compressive strength other mixture components yes
coarse aggregate compressive strength other mixture components yes
fine aggregate compressive strength other mixture components yes
age compressive strength other mixture components yes
alcoholic comsumption mean corpuscular volume genes yes
alcoholic comsumption alkphos genes yes
alcoholic comsumption sgpt genes yes
alcoholic comsumption sgot genes yes
alcoholic comsumption gammagt genes yes
age body mass index (weight in kg/(height in m)2) genes yes
age 2-Hour serum insulin (mu U/ml) genes yes
age diastolic blood pressure (mm Hg) genes yes
age Plasma glucose concentration a 2 hours in an oral glucose tolerance test genes yes
days of the year mean daily temperature of Furtwangen location timeseries
temperature: year 2000, day 50 temperature: year 2000, day 51 location yes
temperature: year 2000, day 50 temperature: year 2000, day 51 location yes
temperature: year 2000, day 50 temperature: year 2000, day 51 location yes
temperature: year 2000, day 50 temperature: year 2000, day 51 location yes
weekend? (binary data) number of cars per 24h at different counting stations in Oberschwaben, Germany weather yes
Outdoor temperature Indoor temperature timeseries yes
Temperature (degree celsius) Ozone (microgram / cubic meter) Car gases emission yes
Temperature (degree celsius) Ozone (microgram / cubic meter) Car gases emission yes
Temperature (degrees Celsius), Davos-See, Switzerland Ozone (microgram / cubic meter), Davos-See, Switzerland Car gases emission yes
day 50: 4 metereological variables day 51: 4 metereological variables location yes
(wind speed ($m/s$), global radiation ($W/m^{2}$), temperature) Ozon concentration ($microgramm/m^{3}$) Car gases emission yes
(displacement, horsepower, weight) (mpg, acceleration) aerodynamics yes
Temperature (degree celsius) Ozone (microgram / cubic meter) Car gases emission yes
latitude of the country’s capital life expectancy at birth for different countries, female, 2000-2005 longitude yes
latitude of the country’s capital life expectancy at birth for different countries, female, 1995-2000 longitude yes
latitude of the country’s capital life expectancy at birth for different countries, female, 1990-1995 longitude yes
latitude of the country’s capital life expectancy at birth for different countries, female, 1985-1990 longitude yes
latitude of the country’s capital life expectancy at birth for different countries, male, 2000-2005 longitude yes
latitude of the country’s capital life expectancy at birth for different countries, male, 1995-2000 longitude yes
latitude of the country’s capital life expectancy at birth for different countries, male, 1990-1995 longitude yes
latitude of the country’s capital life expectancy at birth for different countries, male, 1985-1990 longitude yes
Population with sustainable access to improved drinking water sources (%) total, 2006 Infant mortality rate (per 1 000 live births) both sexes, 2006 health car access yes
stock returns of Hang Seng Bank (0011.HK) stock return of HSBC Hldgs (0005.HK) other HSBC holdings yes
stock returns of Hutchison (0013.HK) stock return of Cheung kong (0001.HK) other Cheung kong holdings yes
stock returns of Cheung Kong (0001.HK) stock return of Sun Hung Kai Prop. (0016.HK) other Sun Hung Kai Prop. holdings yes
open http connections during that minute bytes sent at minute t user type yes
outside temperature in degrees Celsius inside room temperature in degrees Celsius timeseries
par (between 0 and 14, 0 >-> very female, 14 >-> very male) sex guess (0: female or 1: male, the subject’s guess) design choices yes
(Temperature of patient 35C-42C , Occurrence of nausea, Lumbar pain, Urine pushing (continuous need for urination), Micturition pains, Burning of urethra, itch, swelling of urethra outlet) (decision: Inflammation of urinary bladder, decision: Nephritis of renal pelvis origin) age, doctor yes
sunspot area global mean temperature anomalies (deviations from 1961-1990) in °C location yes
Energy use (kg of oil equivalent per capita) for different countries in different years CO2 emissions for different countries in different years country development level yes
GNI (Gross national income) per capita for different countries (in US\$) life expectancy at birth for different countries health car access yes
GNI (Gross national income) per capita for different countries (in US\$) under 5 mortality rate for different countries (deaths per 1000 live births) health car access yes
the average annual rate of change of population the average annual rate of change of total dietary consumption for total population (kcal/day) rate of change of individual dietary consumption yes
the solar radiation in W/m2 the daily average temperature of the air measured at the same location and the same days latitude timeseries
PPFD (Photosynthetic Photon Flux Density) NEP (Net Ecosystem Productivity) Gross Primary Productivity (GPP) yes
PPFDdif (Photosynthetic Photon Flux Density, diffusive) NEP (Net Ecosystem Productivity) PPFDdir (Photosynthetic Photon Flux Density, direct) yes
PPFDdir (Photosynthetic Photon Flux Density, direct) NEP (Net Ecosystem Productivity) PPFDdif (Photosynthetic Photon Flux Density, diffusive) yes
Temperature in degree Celsius CO2 flux at night timeseries
Temperature in degree Celsius CO2 flux at night timeseries
Temperature in degree Celsius CO2 flux at night timeseries
the natural logarithm of the corresponding population the natural logarithm of employment in 1980 in 3102 counties in US wealth yes
time to take weekly measurements (from 1 to 14) protein content of the milk produced by each cow at time X genes yes
size in m2 of appartment/room monthly rent in EUR location yes
MeanTemp (deg Celsius) TotalSnow (cm) precipitation yes
age Relative Spinal bone mineral density genes yes
Mass loss OCTOBER 2012 in % Mass loss APRIL 2012 in % ecosystem yes
Mass loss OCTOBER 2012 in % Mass loss APRIL 2012 in % ecosystem yes
Clay content in soil (in gram per g/kg) Soil moisture at 10cm depth (in %) precipitation yes
Organic C content in soil (in g Carbon/kg) Clay content (in g/kg) concentration of other atoms yes
average precipitation over 1948 to 2004 in mm/day average runoff in over 1948 to 2004 mm/day drainage system quality yes
hour of the day temperature in degree celsius location timeseries
hour of the day load: the total electricity consumption in a region of Turkey in ”MWh” location timeseries
temperature in degree celsius load: the total electricity consumption in a region of Turkey in ”MWh” location timeseries
initial speed of a ball on a ball track for children final speed of a ball on a ball track for children none no
initial speed of a ball on a ball track for children final speed of a ball on a ball track for children none no
social-economic status of pupil’s family language test score IQ yes
cycle time in nanoseconds published performance on a benchmark mix relative to an IBM 370/158-3 energy consumption yes
grey value of a pixel that is chosen randomly from a fixed image. The grey value light intensity seen by a photo diode placed several centimeters away from the screen. none no
position on the ball track where the ball starts time interval between passing the first and the second light barrier none no
position on the ball track where the ball starts time interval between passing the third and the fourth light barrier none no
time interval between passing the third and the fourth light barrier time interval between passing the third and the fourth light barrier none no
pixel vector of grey values of the patch light intensity seen by a photo diode placed several centimeters away from the screen none no
electric voltage time required for passing one round none no
contrast answer correct or not observer yes
time for 1/6 rotation temperature in Degree Celsius none no
Appendix B Synthetic Dataset Generation
B.1 Functions Specifications
Throughout this appendix, $x\in[0,1]$ denotes the min–max rescaled output produced by Algorithm 1. All latent parameters (collectively denoted $\psi$) are rescaled draws of a common latent variable sampled from one of the three base priors (Uniform, Normal, Rayleigh) (see Subsection 4.2). Additionally, since the latent variables were designed to interact with the input in complex ways, the common approach was to choose function hyperparameters that change the shape of the function. However, there still exists an option to add multiplicative noise, which will not be considered here. Lastly, in functions whose design is monotone for simplicity, a Bernoulli fair coin is tossed once per example to decide whether to flip the sign of $y$ (and, for some mechanisms, of $x$ as well) so that both monotone orientations occur with equal frequency. Now, in greater detail, the eight implemented causal functions are:
  1. 1.
    Exponential ($f_{\mbox{exp}}$): this implementation starts by sampling the latent variable rescaling parameters:
    $\mu_{\psi}\sim F_{\mu}(\mu_{\mu},\sigma_{\mu});\sigma_{\psi}\sim F_{\sigma}(\mu_{\sigma},\sigma_{\sigma})$ (10)
    where the choices for $\mu_{\mu},\sigma_{\mu},\mu_{\sigma},\sigma_{\sigma}$ and the functions $F_{\mu},F_{\sigma}$ constitute hyperparameters. Consequently, $y$ can be computed as
    $y_{i}=e^{x_{i}a_{i}},a=\mbox{Rescale}(\psi;\mu_{\psi},\sigma_{\psi}).$ (11)
  2. 2.
    Logarithmic ($f_{\log}$): the implementation of the logarithmic function closely resembles that of the exponential function above; the only change lies in the final function itself, which is
    $y_{i}=\log(x_{i}+a_{i}).$ (12)
  3. 3.
    Inversely proportional ($f_{\mbox{inv}}$): this function also mimics the exponential function in terms of obtaining the rescaled version of the latent variable. Its final expression is
    $y_{i}=\frac{1}{x_{i}+a_{i}}.$ (13)
  4. 4.
    Power law ($f_{\mathrm{pow}}$): the power law function is the last one that resembles the exponential in its architecture. Similarly, the expression used to compute $y$ from the rescaled $\psi$ distribution and $x$ is
    $y_{i}=x_{i}^{a_{i}}.$ (14)
  5. 5.
    Linear ($f_{\mbox{lin}}$): since $f$ must be strictly increasing in $\psi$ (in accordance with our design choice), in order to generate samples using linear functions it suffices to define the distribution at the two endpoints (0 and $1$ respectively)
    $\begin{gathered}\sigma_{0},\sigma_{1}\sim\mbox{Inv-Gamma}\bigl(\gamma_{m},\,\gamma_{v}\bigr)\\[6.0pt] \begin{aligned} \mu_{0}&=0\qquad&\mu_{1}&=\tan\phi\qquad&\phi&\sim\mathcal{U}(0,2\pi)\end{aligned}\end{gathered}$ (15)
    Consequently, the $\psi$ distribution is rescaled according to the sampled values above
    $a=\mbox{Rescale}(\psi;\mu_{0},\sigma_{0})\qquad b=\mbox{Rescale}(\psi;\mu_{1},\sigma_{1}),$ (16)
    where “Rescale” denotes the operation that adjusts the $\psi$ distribution to have mean $\mu$ and standard deviation $\sigma$, while preserving its original shape and point‐mass locations. Having the distributions at the two endpoints ($a$ and $b$), it is possible to obtain $y_{i}$
    $y_{i}=a_{i}+(b_{i}-a_{i})x_{i}.$ (17)
  6. 6.
    Piecewise linear ($f_{\mbox{mix}}$): the piecewise linear function can be implemented based on the linear function with two small adaptations:
    • There is an extra hyperparameter: $\mbox{max}_{K}$, which denotes the maximum number of slices the piecewise model should have; then, $K\sim\mathcal{U}(2,\mbox{max}_{K})$.
    • The smaller size of each slice (when compared to the entire $[0,1]$ interval) requires adapting the rescaling average $\mu$ to ensure the function is continuous:
      $\mu_{k+1}=\mu_{k}+\tan(\phi)(x_{k+1}-x_{k}),$ (18)
      where both $k$ and $k+1$ define the endpoints of each slice.
  7. 7.
    Brownian‐like motion ($f_{\mbox{brown}}$): the Brownian-like function can be seen has a piecewise linear function with two different adaptations:
    • Each slice only contains one point. Thus, there are $N_{S}$ slices.
    • In order to avoid constant oscillation, an extra momentum term is added to the computation of the rescaling variables to ensure the function is smooth and differentiable,
      $\mu_{k+1}\sim N\left(\mu_{k}+\frac{d\mu}{dt}\Delta_{t},\sqrt{\frac{\Delta_{t}^{P}}{P}}\right),$ (19)
      where $\frac{d\mu}{dt}$ is computed using a polynomial approximation, $P$ is a hyperparameter and the expression for the standard deviation is obtained from that of integrated Brownian motion.
  8. 8.
    Polynomial ($f_{\mbox{poly}}$): finally, the polynomial function starts by sampling the order of the polynomial: $o\sim\mathcal{U}(2,\mbox{max}_{o})$, where the maximum value ($\mbox{max}_{o}$) is an hyperparameter. Afterwards, $o+1$ random points are selected from $x$, which will represent our slices. Then, similarly to the piecewise linear function, we obtain the rescaled $\psi$ at each slice. Finally, for each $y_{i}$, given the corresponding resampled $\psi$,
    $\psi_{1_{i}},\psi_{k_{i}},\dots,\psi_{o_{i}}.$ (20)
    A polynomial of order $o$ is then fitted to the following pairs of points:
    $p_{i}=\mbox{fit}(x_{[k]},\psi_{[k]_{i}}).$ (21)
    Finally, $y_{i}$ can be computed as:
    $y_{i}=p_{i}(0)+p_{i}(1)x+p_{i}(2)x^{2}+...p_{i}(o)x^{o}.$ (22)
B.2 Examples from the Synthetic Dataset
Figure 4: This Figure displays 32 examples randomly sampled from the developed synthetic dataset.
Appendix C SynthNN specifications
In this appendix, we provide the full technical description of the convolutional neural network (CNN) used for binary classification, the image‐conversion and labeling procedures, and relevant implementation details omitted from the main text.
C.1 Data preprocessing
The data is originally presented as two paired vectors ($x$ and $y$). This is converted into an image using the following steps:
  1. 1.
    Min–maxscale both $\mathbf{x}$ and $\mathbf{y}$ to $[0,1]$.
  2. 2.
    Convert scaled coordinates to integer pixel indices:
    $i=\lfloor x\,(N-1)\rfloor,\quad j=\lfloor y\,(N-1)\rfloor.$
  3. 3.
    Initialize an $N\times N$ zero matrix and set $I[j,i]=1$ for each point $(i,j)$, where $N=50$.
Afterwards, since there is no inherent true direction, half the images must be assigned to the $Y\xrightarrow{}X$ direction, and their axis must be switched. This is required to ensure the network sees $50\%$ of $X\xrightarrow{}Y$ and $50\%$ of $Y\xrightarrow{}X$, so that it has no prior bias. Since the examples are drawn from different combinations of design choices (see Subsection 4.3), the examples within each set of design choices were guaranteed to have half of each class, to ensure the network does not learn any biases towards the design choices (and not the causal direction). Finally, in order to improve the results, a Gaussian filter is applied to the images with $\sigma=0.5$, exclusively during training.
C.2 Architecture
The neural network has a total of 1,739,777 trainable parameters, containing:
  • Convolutional blocks: three 2D convolutional layers (3×3 kernels), each followed by a 2×2 max‐pooling, with filter counts doubling each block (32, 64, 128).
  • Dense layers: three fully connected layers of sizes 256, 128, and 64, respectively.
  • Activation: ReLU for all hidden layers, sigmoid activation for the final output neuron.
  • Regularization: $\ell_{2}$ weight decay with coefficient $\lambda=0.01$ on all kernels.
  • Loss function: binary cross‐entropy.
  • Optimizer: Adam $\bigl(\alpha=10^{-4}\bigr)$.
The sequential architecture of SynthNN can be seen in Table 5.
Layer (type) Output Shape # Parameters
Conv2D (3×3, 32 filters) (None, 50, 50, 32) 320
MaxPooling2D (2×2) (None, 25, 25, 32) 0
Conv2D (3×3, 64 filters) (None, 25, 25, 64) 18,496
MaxPooling2D (2×2) (None, 13, 13, 64) 0
Conv2D (3×3, 128 filters) (None, 13, 13, 128) 73,856
MaxPooling2D (2×2) (None, 7, 7, 128) 0
Flatten (None, 6 272) 0
Dense (256 units) (None, 256) 1,605,888
Dense (128 units) (None, 128) 32,896
Dense (64 units) (None, 64) 8,256
Dense (1 unit, sigmoid) (None, 1) 65
Table 5: Detailed layer-by-layer summary of SynthNN
C.3 Evaluation on the Tübingen dataset
The output of the neural network represents the estimated posterior probability that the input data has the causal direction $X\xrightarrow{}Y$. However, in order to improve prediction consistency at test time, the following procedure is implemented:
  1. 1.
    Remove outliers beyond the $90\%$ quantile.
  2. 2.
    Generate two images per pair: one for $(X,Y)$ and one for $(Y,X)$.
  3. 3.
    Obtain model predictions $\hat{p}_{X\xrightarrow{}Y}$, $\hat{p}_{Y\xrightarrow{}X}$.
  4. 4.
    Output the asymmetry score:
    $s=\frac{\hat{p}_{X\xrightarrow{}Y}-\hat{p}_{Y\xrightarrow{}X}}{\hat{p}_{X\xrightarrow{}Y}+\hat{p}_{Y\xrightarrow{}X}}.$